# These $\altaffiliationThese authors contributed equally to this work.$

### v1

@@ 4 Basis set convergence for excitation energies of charged and neutral excitations
As discussed in the last Sec. 3, an essential requirement for basis sets used in excited-state electronic structure methods is numerical stability, ensured for example by keeping the condition number of the overlap matrix $\mathbf{S}$ sufficiently low. Equally important is that the basis sets enable fast convergence of excited-state energies with respect to the basis set size. In this section, we analyze the convergence behavior of two types of excitations: (i) charged excitations, corresponding to quasiparticle (QP) energies as obtained from $GW$, and (ii) charge-neutral excitations, such as those calculated from BSE or TDDFT 13.
@@
Charged excitations: GW quasiparticle energies. It is well established that absolute $GW$ quasiparticle energies $\varepsilon_{n}$ converge slowly with increasing basis set size $N_{\text{bf}}$ 44, 45. However, energy differences between QP states, such as the HOMO-LUMO gap, converge much faster 45, 26. We rationalize this behavior in this section, starting by dividing the QP energies into ionization potentials (IPs) and electron affinities (EAs) 13:
@@
Here, $E_{N\pm 1}^{n}$ is the $n$th excited state energy of the $N{\pm}1$ electron system and $\varepsilon_{\text{F}}$ is the Fermi energy. .
According to the second Hohenberg-Kohn theorem 46, the ground state energy $E_{N}^{0}$ of KS-DFT converges from above towards the complete-basis-set (CBS) limit as the basis size increases, i.e.,
@@
We make this assumption also for the excited-state energies $E_{N}^{n}$ and $E_{N\pm 1}^{n}$, following the Hylleraas-Undheim-MacDonald theorem 47, 48:
@@
Summarizing, we can expect relatively fast convergence of BSE excitation energies and $GW$ QP energy differences like the $GW$ HOMO-LUMO gap with the basis set size, given that the basis sets are well optimized, while absolute values of $GW$ QP energy levels are hard to converge with the basis set size. For crystals, relative QP energy differences are often of central interest, like the direct or indirect band gap, such that we focus on relative QP energy differences in this work.
5 Basis-Set Generation Recipe
A Gaussian basis set (8) for an element $A$ is constructed by specifying the total number of basis functions $N_{\text{bf}}^{A}$, selecting the number of functions $\phi_{\mu}(\mathbf{r})$ for each angular momentum quantum number $l$. Each $\phi_{\mu}(\mathbf{r})$ consists of a linear combination (7) of primitive Gaussians characterized by exponential decay parameters $\beta_{\mu,i}$ and contraction coefficients $\alpha_{\mu,i}$. One motivation for this contraction scheme is to better approximate Slater-type orbitals, which decay exponentially as $\exp(-\zeta r)$ and represent the shape of atomic orbitals in the hydrogen-like model 31. The parameters of each contracted function $\phi_{\mu}(\mathbf{r})$ are then optimized to reproduce one or more atomic or molecular properties. These may include, for instance, the correlation energy of neutral atoms 42 or negatively charged atoms 34 or the ground state energy molecules from DFT with the PBE functional 35.
By increasing the number of basis functions $N_{\text{bf}}^{A}$ of the basis set, one can construct hierarchical families of basis sets with systematic improvements. The cc-pVXZ basis set family, X = D, T, Q, $\ldots$ 42 is designed to systematically improve the correlation energy of molecules obtained from post-Hartree-Fock methods like MP2 40, the random phase approximation 49 or coupled cluster 40. The aug-cc-pVXZ basis set family, X = D, T, Q, $\ldots$ 34 is constructed to yield systematic improvements of electron affinities of molecules computed from post-Hartree-Fock methods. In contrast, the XZVP-MOLOPT basis set family, X = S, D, T, Q 35, 37, 38 is designed to obtain basis-set converged groundstate DFT calculations of large molecules, crystals, liquids and material interfaces. For a more complete review of Gaussian basis sets, we refer to Ref. 40.
@@ 5 Basis-Set Generation Recipe
H, He aug-$c$-pVDZ 3s, 2p 9 cc-pVDZ 42, 43 2s, 1p 1s, 1p 0.030 (H), 0.072 (He)
Li – Ne aug-$c$-pVDZ 4s, 3p, 2d 23 cc-pVDZ 42, 50 3s, 2p, 1d 1s, 1p, 1d 0.006 (Li) – 0.106 (Ne)
Na – Cl aug-$c$-pVDZ 5s, 4p, 2d 27 cc-pVDZ 51, 50 4s, 3p, 1d 1s, 1p, 1d 0.006 (Na) – 0.047 (Cl)
H, He aug-$c$-pVTZ 4s, 3p, 2d 23 cc-pVTZ 42, 43 3s, 2p, 1d 1s, 1p, 1d 0.025 (H), 0.052 (He)
Li – Ne aug-$c$-pVTZ 5s, 4p, 3d, 2f 46 cc-pVTZ 42, 50 4s, 3p, 2d, 1f 1s, 1p, 1d, 1f 0.008 (Li) – 0.092 (Ne)
Na – Cl aug-$c$-pVTZ 6s, 5p, 3d, 2f 50 cc-pVTZ 51, 50 5s, 4p, 2d, 1f 1s, 1p, 1d, 1f 0.007 (Na) – 0.042 (Cl)
H, He aug-$c$-pVQZ 5s, 4p, 3d, 2f 46 cc-pVQZ 42, 43 4s, 3p, 2d, 1f 1s, 1p, 1d, 1f 0.024 (H), 0.048 (He)
Li – Ne aug-$c$-pVQZ 6s, 5p, 4d, 3f, 2g 80 cc-pVQZ 42, 50 5s, 4p, 3d, 2f, 1g 1s, 1p, 1d, 1f, 1g 0.006 (Li) – 0.082 (Ne)
Na – Cl aug-$c$-pVQZ 7s, 6p, 4d, 3f, 2g 84 cc-pVQZ 50, 51 6s, 5p, 3d, 2f, 1g 1s, 1p, 1d, 1f, 1g 0.005 (Na) – 0.038 (Cl)
H, He aug-SZV-M.-ae-mini 3s, 1p 6 STO-3G 31 1s 2s, 1p 0.065 (H), 0.090 (He)
Li – Ne aug-SZV-M.-ae-mini 3s, 2p 9 STO-3G 31, 52 2s, 1p 1s, 1p 0.048 (Li) – 0.200 (Ne)
Na – Cl aug-SZV-M.-ae-mini 4s, 3p 13 STO-3G 52 3s, 2p 1s, 1p 0.050 (Na) – 0.080 (Cl)
H aug-SZV-M.-ae-SR 3s, 1p 6 STO-3G 31 1s 2s, 1p 0.065 (H)
C, N, O aug-SZV-M.-ae-SR 3s, 2p, 1d 14 STO-3G 31 2s, 1p 1s, 1p, 1d 0.115 (C) – 0.162 (O)
H, He aug-SZV-M.-ae 3s, 1p 6 STO-6G 31 1s 2s, 1p 0.065 (H), 0.090 (He)
Li – Ne aug-SZV-M.-ae 3s, 2p, 1d 14 STO-6G 31, 52 2s, 1p 1s, 1p, 1d 0.031 (Li) – 0.200 (Ne)
Na – Cl aug-SZV-M.-ae 4s, 3p, 1d 18 STO-6G 52 3s, 2p 1s, 1p, 1d 0.050 (Na) – 0.077 (Cl)
H, He aug-DZVP-M.-ae 3s, 2p 9 SVP-M.-ae 37, 38 2s, 1p 1s, 1p 0.035 (H), 0.060 (He)
Li – Ne aug-DZVP-M.-ae 4s, 3p, 2d, 1f 30 SVP-M.-ae 37, 38 3s, 2p, 1d 1s, 1p, 1d, 1f 0.025 (Li) – 0.100 (Ne)
Na – Cl aug-DZVP-M.-ae 5s, 4p, 2d, 1f 34 SVP-M.-ae 37, 38 4s, 3p, 1d 1s, 1p, 1d, 1f 0.045 (Na) – 0.077 (Cl)
H, He aug-TZVP-M.-ae 4s, 3p, 2d 23 TZVPP-M.-ae 37, 38 3s, 2p, 1d 1s, 1p, 1d 0.030 (H), 0.050 (He)
Li – Ne aug-TZVP-M.-ae 6s, 4p, 3d, 2f, 1g 56 TZVPP-M.-ae 37, 38 5s, 2p, 2d, 1f 1s, 1p, 1d, 1f, 1g 0.025 (Li) – 0.100 (Ne)
Na – Cl aug-TZVP-M.-ae 6s, 6p, 4d, 2f, 1g 67 TZVPP-M.-ae 37, 38 5s, 5p, 3d, 1f 1s, 1p, 1d, 1f, 1g 0.025 (Na) – 0.130 (Cl)
@@
We recall the strategy behind the construction of Dunning’s aug-cc-pVXZ basis sets 34, 53 as summarized in Table 1, to motivate our approach of constructing basis sets. The aug-cc-pVXZ basis sets are designed to compute electron affinities, which requires accurate description of the lowest unoccupied molecular orbital (LUMO) in KS-DFT. The LUMO is typically much more diffuse than occupied orbitals as indicated by the decay length $\zeta_{n}\hskip 1.69998pt{\approx}\hskip 1.69998pt\hbar/{\sqrt{2m|\varepsilon_{n}|}}$ of a molecular orbital outside of the molecule; $\varepsilon_{n}$ is the orbital energy of a bound state and $m$ the electron mass. For occupied orbitals, we usually have $|\varepsilon_{n}|\hskip 1.69998pt{<}\hskip 1.69998pt5$ eV while $|\varepsilon_{\text{LUMO}}|\hskip 1.69998pt{\approx}\hskip 1.69998pt0$ eV leads to a long decay length $\zeta_{\text{LUMO}}$, indicating a diffuse LUMO. Standard cc-pVXZ basis sets lack the necessary diffuse functions to describe such orbitals. To address this, aug-cc-pVXZ adds an uncontracted diffuse Gaussian $\phi_{\mu}(\mathbf{r})\hskip 1.69998pt{=}\hskip 1.69998ptY_{l}^{m}(\theta_{A},\varphi_{A})\,r_{A}^{l}\,\alpha_{l}\,\exp\left(-\beta_{l}\,r_{A}^{2}\right)$ for each angular momentum $l$ present in cc-pVXZ. The exponent $\beta_{l}$ is optimized to match the correlation energy of the corresponding anion at the complete-basis-set limit. These augmented basis sets have proven effective for excited-state properties, such as $GW$ HOMO-LUMO gaps and TDDFT or GWGW+BSE excitation energies 54, 45, 55.
@@
While aug-cc-pVXZ basis sets yield good accuracy for excitation energies of small and medium-sized molecules, their application to large or periodic systems is severely limited by numerical issues. The inclusion of very diffuse Gaussians results in a large condition number of the overlap matrix $\mathbf{S}$, making the computation (6) of $\mathbf{S}^{-1/2}$ numerically unstable, as discussed in Sec. 3. This instability often leads to convergence problems of the self-consistent-field cycle when treating solids or large molecules. Consequently, the use of aug-cc-pVXZ is usually prohibitive for solids and large molecules.
@@
We develop a family of augmented all-electron (ae) MOLOPT basis sets specifically targeted for excited-state calculations of large molecules and solids, following the analysis of numerical stability from Sec. 3 and basis set convergence from Sec. 4. We apply an augmentation strategy to the basis sets STO-3G 31, STO-6G 31, SVP-MOLOPT-ae 37, 38, and TZVPP-MOLOPT-ae 37, 38, resulting in the aug-SZV-MOLOPT-mini-ae, aug-SZV-MOLOPT-ae-SR, aug-SZV-MOLOPT-ae, aug-DZVP-MOLOPT-ae, and aug-TZVP-MOLOPT-ae basis sets, respectively, see Table 1. We add one additional angular momentum shell beyond $l_{\text{max}}$ of the root basis, and introduce one new diffuse Gaussian primitive with a smaller exponent per angular momentum to improve radial flexibility. We optimize the added basis functions to reproduce the lowest five $G_{0}W_{0}$-BSE@PBE0 excitation energies calculated with the aug-cc-pV5Z basis set for Thiel’s set 56 (for elements H, C, N, O) and the available molecular set from Ref. 35 for the other elements up to chlorine. In the objective function, we also ensure that the condition number of the basis set stays limited, as in previous optimizations of Gaussian basis sets 35, 36.
@@
Figure 2: aug-SZV-MOLOPT-ae-mini, aug-SZV-MOLOPT-ae-SR, aug-SZV-MOLOPT-ae and aug-DZVP-MOLOPT-ae basis sets developed in this work for carbon in CP2K basis set format 57, 53. Numbers marked in red have been optimized to match BSE excitation energies of the molecules contained in Thiel’s set. The black numbers are the parameters taken from STO-3G 31 (for aug-SZV-MOLOPT-ae-SR), STO-6G 31 (for aug-SZV-MOLOPT-ae) and from SVP-MOLOPT-PBE-ae 37, 38 (for aug-DZVP-MOLOPT-ae). The basis sets for other atoms and all corresponding auxiliary RI basis sets are listed in the supporting information S3, S4.
@@
where $\alpha_{i,j}$ can be found in the literature 31, 53 and the exponents are
@@
6 auxiliary RI basis Set Generation
The resolution-of-the-identity (RI) technique is widely used to reduce the computational cost of quantum chemical methods 58. In RI, four-center integrals

### v2

@@ 2.3 Basis set convergence for excitation energies of charged and neutral excitations
As discussed in the last Sec. 2.2, an essential requirement for basis sets used in excited-state electronic structure methods is numerical stability, ensured for example by keeping the condition number of the overlap matrix $\mathbf{S}$ sufficiently low. Equally important is that the basis sets enable fast convergence of excited-state energies with respect to the basis set size. In this section, we analyze the convergence behavior of two types of excitations: (i) charged excitations, corresponding to quasiparticle (QP) energies as obtained from $GW$, and (ii) charge-neutral excitations, such as those calculated from BSE or TDDFT Golze2019.
@@
Charged excitations: GW quasiparticle energies. It is well established that absolute $GW$ quasiparticle energies $\varepsilon_{n}$ converge slowly with increasing basis set size $N_{\text{bf}}$ gw100, wilhelm2016 and thus require basis set extrapolation Bruneval2020, Baum2025. However, energy differences between QP states, such as the HOMO-LUMO gap, converge much faster wilhelm2016, foerster2022quasiparticle. We rationalize this behavior in this section, starting by dividing the QP energies into ionization potentials (IPs) and electron affinities (EAs) Golze2019:
@@
Here, $E_{N\pm 1}^{n}$ is the $n$th excited state energy of the $N{\pm}1$ electron system and $\varepsilon_{\text{F}}$ is the Fermi energy.
According to the second Hohenberg-Kohn theorem Hohenberg1964, the ground state energy $E_{N}^{0}$ of KS-DFT converges from above toward the complete-basis-set (CBS) limit as the basis size increases, i.e.,
@@
We make this assumption also for the excited-state energies $E_{N}^{n}$ and $E_{N\pm 1}^{n}$, following the Hylleraas-Undheim-MacDonald theorem Hylleraas1930, MacDonald1933:
@@
Summarizing, we can expect relatively fast convergence of BSE excitation energies and $GW$ QP energy differences like the $GW$ HOMO-LUMO gap with the basis set size, given that the basis sets are well optimized, while absolute values of $GW$ QP energy levels are hard to converge with the basis set size.
2.4 Basis set generation recipe
A Gaussian basis set (8) for an element $A$ is constructed by specifying the total number of basis functions $N_{\text{bf}}^{A}$, selecting the number of functions $\phi_{\mu}(\mathbf{r})$ for each angular momentum quantum number $l$. Each $\phi_{\mu}(\mathbf{r})$ consists of a linear combination (7) of primitive Gaussians characterized by exponential decay parameters $\beta_{\mu,i}$ and contraction coefficients $\alpha_{\mu,i}$. One motivation for this contraction scheme is to better approximate Slater-type orbitals, which decay exponentially as $\exp(-\zeta r)$ and represent the shape of atomic orbitals in the hydrogen-like model hehre1969a. The parameters of each contracted function $\phi_{\mu}(\mathbf{r})$ are then optimized to reproduce one or more atomic or molecular properties. These may include, for instance, the correlation energy of neutral atoms dunning1989a or negatively charged atoms kendall1992a or the ground state energy molecules from DFT with the PBE functional Vandevondele2007.
By increasing the number of basis functions $N_{\text{bf}}^{A}$ of the basis set, one can construct hierarchical families of basis sets with systematic improvements. The cc-pVXZ basis set family, X = D, T, Q, $\ldots$ dunning1989a is designed to systematically improve the correlation energy of molecules obtained from post-Hartree-Fock methods like MP2 jensen, the random phase approximation Ren2012review or coupled cluster jensen. The aug-cc-pVXZ basis set family, X = D, T, Q, $\ldots$ kendall1992a is constructed to yield systematic improvements of electron affinities of molecules computed from post-Hartree-Fock methods. In contrast, the XZVP-MOLOPT basis set family, X = S, D, T, Q Vandevondele2007, MOLOPT_PBE_ae, PhDTMueller is designed to obtain basis-set converged groundstate DFT calculations of large molecules, crystals, liquids and material interfaces. For a more complete review of Gaussian basis sets, we refer to Ref. jensen.
@@ 2.4 Basis set generation recipe
H, He aug-$c$-pVDZ 3s, 2p 9 cc-pVDZ dunning1989a, woon1994a 2s, 1p 1s, 1p 0.030 (H), 0.072 (He)
Li – Ne aug-$c$-pVDZ 4s, 3p, 2d 23 cc-pVDZ dunning1989a, prascher2011a 3s, 2p, 1d 1s, 1p, 1d 0.006 (Li) – 0.106 (Ne)
Na – Cl aug-$c$-pVDZ 5s, 4p, 2d 27 cc-pVDZ woon1993a, prascher2011a 4s, 3p, 1d 1s, 1p, 1d 0.006 (Na) – 0.047 (Cl)
H, He aug-$c$-pVTZ 4s, 3p, 2d 23 cc-pVTZ dunning1989a, woon1994a 3s, 2p, 1d 1s, 1p, 1d 0.025 (H), 0.052 (He)
Li – Ne aug-$c$-pVTZ 5s, 4p, 3d, 2f 46 cc-pVTZ dunning1989a, prascher2011a 4s, 3p, 2d, 1f 1s, 1p, 1d, 1f 0.008 (Li) – 0.092 (Ne)
Na – Cl aug-$c$-pVTZ 6s, 5p, 3d, 2f 50 cc-pVTZ woon1993a, prascher2011a 5s, 4p, 2d, 1f 1s, 1p, 1d, 1f 0.007 (Na) – 0.042 (Cl)
H, He aug-$c$-pVQZ 5s, 4p, 3d, 2f 46 cc-pVQZ dunning1989a, woon1994a 4s, 3p, 2d, 1f 1s, 1p, 1d, 1f 0.024 (H), 0.048 (He)
Li – Ne aug-$c$-pVQZ 6s, 5p, 4d, 3f, 2g 80 cc-pVQZ dunning1989a, prascher2011a 5s, 4p, 3d, 2f, 1g 1s, 1p, 1d, 1f, 1g 0.006 (Li) – 0.082 (Ne)
Na – Cl aug-$c$-pVQZ 7s, 6p, 4d, 3f, 2g 84 cc-pVQZ prascher2011a, woon1993a 6s, 5p, 3d, 2f, 1g 1s, 1p, 1d, 1f, 1g 0.005 (Na) – 0.038 (Cl)
H, He aug-SZV-M.-ae-mini 3s, 1p 6 STO-3G hehre1969a 1s 2s, 1p 0.065 (H), 0.090 (He)
Li – Ne aug-SZV-M.-ae-mini 3s, 2p 9 STO-3G hehre1969a, hehre1970a 2s, 1p 1s, 1p 0.048 (Li) – 0.200 (Ne)
Na – Cl aug-SZV-M.-ae-mini 4s, 3p 13 STO-3G hehre1970a 3s, 2p 1s, 1p 0.050 (Na) – 0.080 (Cl)
H aug-SZV-M.-ae-SR 3s, 1p 6 STO-3G hehre1969a 1s 2s, 1p 0.065 (H)
C, N, O aug-SZV-M.-ae-SR 3s, 2p, 1d 14 STO-3G hehre1969a 2s, 1p 1s, 1p, 1d 0.115 (C) – 0.162 (O)
H, He aug-SZV-M.-ae 3s, 1p 6 STO-6G hehre1969a 1s 2s, 1p 0.065 (H), 0.090 (He)
Li – Ne aug-SZV-M.-ae 3s, 2p, 1d 14 STO-6G hehre1969a, hehre1970a 2s, 1p 1s, 1p, 1d 0.031 (Li) – 0.200 (Ne)
Na – Cl aug-SZV-M.-ae 4s, 3p, 1d 18 STO-6G hehre1970a 3s, 2p 1s, 1p, 1d 0.050 (Na) – 0.077 (Cl)
H, He aug-DZVP-M.-ae 3s, 2p 9 SVP-M.-ae MOLOPT_PBE_ae, PhDTMueller 2s, 1p 1s, 1p 0.035 (H), 0.060 (He)
Li – Ne aug-DZVP-M.-ae 4s, 3p, 2d, 1f 30 SVP-M.-ae MOLOPT_PBE_ae, PhDTMueller 3s, 2p, 1d 1s, 1p, 1d, 1f 0.025 (Li) – 0.100 (Ne)
Na – Cl aug-DZVP-M.-ae 5s, 4p, 2d, 1f 34 SVP-M.-ae MOLOPT_PBE_ae, PhDTMueller 4s, 3p, 1d 1s, 1p, 1d, 1f 0.045 (Na) – 0.077 (Cl)
H, He aug-TZVP-M.-ae 4s, 3p, 2d 23 TZVPP-M.-ae MOLOPT_PBE_ae, PhDTMueller 3s, 2p, 1d 1s, 1p, 1d 0.030 (H), 0.050 (He)
Li – Ne aug-TZVP-M.-ae 6s, 4p, 3d, 2f, 1g 56 TZVPP-M.-ae MOLOPT_PBE_ae, PhDTMueller 5s, 2p, 2d, 1f 1s, 1p, 1d, 1f, 1g 0.025 (Li) – 0.100 (Ne)
Na – Cl aug-TZVP-M.-ae 6s, 6p, 4d, 2f, 1g 67 TZVPP-M.-ae MOLOPT_PBE_ae, PhDTMueller 5s, 5p, 3d, 1f 1s, 1p, 1d, 1f, 1g 0.025 (Na) – 0.130 (Cl)
@@
We recall the strategy behind the construction of Dunning’s aug-cc-pVXZ basis sets kendall1992a, pritchard2019a as summarized in Table 1, to motivate our approach of constructing basis sets. The aug-cc-pVXZ basis sets are designed to compute electron affinities, which requires accurate description of the lowest unoccupied molecular orbital (LUMO) in KS-DFT. The LUMO is typically much more diffuse than occupied orbitals as indicated by the decay length $\zeta_{n}\hskip 1.69998pt{\approx}\hskip 1.69998pt\hbar/{\sqrt{2m|\varepsilon_{n}|}}$ of a molecular orbital outside of the molecule; $\varepsilon_{n}$ is the orbital energy of a bound state and $m$ the electron mass. For occupied orbitals, we usually have $|\varepsilon_{n}|\hskip 1.69998pt{<}\hskip 1.69998pt5$ eV while $|\varepsilon_{\text{LUMO}}|\hskip 1.69998pt{\approx}\hskip 1.69998pt0$ eV leads to a long decay length $\zeta_{\text{LUMO}}$, indicating a diffuse LUMO. Standard cc-pVXZ basis sets lack the necessary diffuse functions to describe such orbitals. To address this, aug-cc-pVXZ adds an uncontracted diffuse Gaussian $\phi_{\mu}(\mathbf{r})\hskip 1.69998pt{=}\hskip 1.69998ptY_{l}^{m}(\theta_{A},\varphi_{A})\,r_{A}^{l}\,\alpha_{l}\,\exp\left(-\beta_{l}\,r_{A}^{2}\right)$ for each angular momentum $l$ present in cc-pVXZ. The exponent $\beta_{l}$ is optimized to match the correlation energy of the corresponding anion at the complete-basis-set limit. These augmented basis sets have proven effective for excited-state properties, such as $GW$ HOMO-LUMO gaps and TDDFT or GWGW+BSE excitation energies Jacquemin2015, wilhelm2016, Knysh2024a.
@@
While aug-cc-pVXZ basis sets yield good accuracy for excitation energies of small and medium-sized molecules, their application to large systems is severely limited by numerical issues. The inclusion of very diffuse Gaussians results in a large condition number of the overlap matrix $\mathbf{S}$, making the computation (6) of $\mathbf{S}^{-1/2}$ numerically unstable, as discussed in Sec. 2.2. This instability often leads to convergence problems of the self-consistent-field cycle when treating large molecules. Consequently, the use of aug-cc-pVXZ is usually prohibitive for large molecules.
@@
We develop a family of augmented all-electron (ae) MOLOPT basis sets specifically targeted for excited-state calculations of large molecules , following the analysis of numerical stability from Sec. 2.2 and basis set convergence from Sec. 2.3. We apply an augmentation strategy to the basis sets STO-3G hehre1969a, STO-6G hehre1969a, SVP-MOLOPT-ae MOLOPT_PBE_ae, PhDTMueller, and TZVPP-MOLOPT-ae MOLOPT_PBE_ae, PhDTMueller, resulting in the aug-SZV-MOLOPT-mini-ae, aug-SZV-MOLOPT-ae-SR, aug-SZV-MOLOPT-ae, aug-DZVP-MOLOPT-ae, and aug-TZVP-MOLOPT-ae basis sets, respectively, see Table 1. The "mini" bases correspond to smaller, more compact versions of the corresponding regular basis sets, and are therefore well-suited for intensive calculations in larger systems with a small cost in accuracy, whereas the "SR" (short-range) basis sets have been generated with less diffuse primitives and are therefore intended to reduce the computational cost in condensed phase systems at a similar accuracy. As such, we expect the following accuracy hierarchy to hold: aug-SZV-MOLOPT-ae-mini < aug-SZV-MOLOPT-ae-SR < aug-DZVP-MOLOPT-ae < aug-TZVP-MOLOPT-ae. We add one additional angular momentum shell beyond $l_{\text{max}}$ of the root basis, and introduce one new diffuse Gaussian primitive with a smaller exponent per angular momentum to improve radial flexibility. One should therefore note that we are using the term augmentation in a broader way than the usual meaning that only involves diffuse functions kendall1992a. The name was chosen to obtain a compact and practical name for the new family of basis sets and underline that these are built on and retain the stability of the MOLOPT basis set, but with better excited-state properties. We optimize the added basis functions to reproduce the lowest five $G_{0}W_{0}$-BSE@PBE0 excitation energies calculated with the aug-cc-pV5Z basis set for Thiel’s set Schreiber2008 (for elements H, C, N, O) and the available molecular set from Ref. Vandevondele2007 for the other elements up to chlorine. In the objective function, we also ensure that the condition number of the basis set stays limited, as in previous optimizations of Gaussian basis sets Vandevondele2007, YeBerkelbach2022. In practice, the optimization is performed using Powell’s algorithm Powell1964, which is a local optimizer and may therefore converge to a local minimum. However, since the optimization in any case depends on the choice of the rather small molecular training set, reaching the global minimum is not a strict requirement to generate high-quality basis sets. Instead, we carefully validate the resulting basis functions on a much larger benchmark set containing 247 molecules, ensuring their general reliability.
@@
Figure 2: aug-SZV-MOLOPT-ae-mini, aug-SZV-MOLOPT-ae-SR, aug-SZV-MOLOPT-ae and aug-DZVP-MOLOPT-ae basis sets developed in this work for carbon in CP2K basis set format cp2k, pritchard2019a. Numbers marked in red have been optimized to match BSE excitation energies of the molecules contained in Thiel’s set. The black numbers are the parameters taken from STO-3G hehre1969a (for aug-SZV-MOLOPT-ae-SR), STO-6G hehre1969a (for aug-SZV-MOLOPT-ae) and from SVP-MOLOPT-PBE-ae MOLOPT_PBE_ae, PhDTMueller (for aug-DZVP-MOLOPT-ae). The basis sets for other atoms and all corresponding auxiliary RI basis sets are listed in the supporting information S3, S4.
@@
where $\alpha_{i,j}$ can be found in the literature hehre1969a, pritchard2019a and the exponents are
@@
[…]
2.5 PBE0 and GW HOMO-LUMO gaps
We compute HOMO-LUMO gaps using PBE0 and $G_{0}W_{0}@$PBE0 for a subset of 247 molecules from the GWGW5000 benchmark set (see App. A for the description of the benchmark set and the computational details). Fig. 3a,b shows the results for the aug-MOLOPT basis sets introduced in this work, together with the aug-cc-pVXZ, cc-pVXZ, and MOLOPT basis sets.
where $N_{\text{mol}}\hskip 1.69998pt{=}\hskip 1.69998pt247$ is the number of molecules and $\Delta_{i}^{B}$ is the PBE0 HOMO-LUMO gap of molecule $i$ computed with basis set $B$. While MAD captures the average accuracy, we also report the 95th percentile error (95PE) to quantify the worst-case deviations of the worst 5% of molecules. The aug-MOLOPT basis sets show systematic improvement from aug-SZV-MOLOPT-ae to aug-TZVP-MOLOPT-ae, with both MAD and 95PE decreasing toward the CBS limit. The aug-TZVP-MOLOPT basis has a MAD of just 14 meV. The aug-cc-pVXZ basis sets generally show even smaller deviations at equivalent basis size—e.g., aug-cc-pVTZ is closer to the CBS than aug-TZVP-MOLOPT-ae. This is expected, as aug-cc-pVXZ are specifically optimized for electron affinities and include very diffuse functions well suited for describing the LUMO. In contrast, the cc-pVXZ and MOLOPT families exhibit significantly larger errors and slower convergence for the HOMO-LUMO gap, reflecting their optimization for ground-state energies rather than excited states. Overall, the aug-MOLOPT basis sets provide fast convergence of HOMO-LUMO gaps, while maintaining a well-conditioned overlap matrix (see Fig. 1).
Fig. 3b shows the basis set convergence of the four basis set families for $G_{0}W_{0}@$PBE0 HOMO-LUMO gaps, using aug-cc-pV5Z as the CBS reference. The aug-MOLOPT basis sets exhibit consistently small deviations from the CBS. Notably, the MAD of the small aug-SZV-MOLOPT-ae basis is 160 meV, better than the larger aug-cc-pVDZ basis, which yields a MAD of 220 meV. Likewise, aug-DZVP-MOLOPT-ae achieves a MAD of just 60 meV, below the 80 meV deviation of the larger aug-cc-pVTZ basis. This finding suggests that the aug-MOLOPT basis sets are the ideal choice for $GW$ HOMO-LUMO gap calculations to ensure both fast basis set convergence of $GW$ HOMO-LUMO gaps and numerical stability for large molecules . Again, the nonaugmented cc-pVXZ and MOLOPT basis sets exhibit larger and more slowly converging errors for $G_{0}W_{0}@$PBE0 HOMO-LUMO gaps. For example, the MAD of the large cc-pV5Z basis is 240 meV—only slightly lower than the minimal aug-SZV-MOLOPT-ae-mini basis which has a MAD of 270 meV. When excluding molecules with diffuse LUMOs (defined as LUMO eigenvalues above – 2 eV), the MAD for cc-pV5Z decreases by an order of magnitude to 30 meV (Fig. S2 in supporting information). This indicates that these basis sets were optimized for ground-state properties, and lacking diffuse functions, they are inadequate for accurately describing $GW$ gaps in systems with unoccupied states that significantly extend into the vacuum.
2.6 GW+BSE and TDDFT excitation energies
[…]
Fig. 3d shows the basis set convergence of the first ten excitation energies computed with TDDFT (PBE0). As before, the aug-MOLOPT basis sets exhibit systematic improvement with increasing basis size. However, in contrast to the BSE case, the aug-cc-pVXZ basis sets outperform the aug-MOLOPT family: for example, the MAD of aug-DZVP-MOLOPT-ae is 100 meV, whereas aug-cc-pVDZ achieves a significantly lower MAD of 17 meV. Comparing with BSE results in Fig. 3c, the aug-cc-pVXZ basis sets converge more rapidly for TDDFT than for BSE, while the aug-MOLOPT sets show similarly fast convergence for both methods. We attribute this difference to the design philosophy behind the basis sets: the aug-MOLOPT sets were specifically optimized for BSE excitation energies (albeit on a different training set, the Thiel’s set Schreiber2008), whereas the aug-cc-pVXZ family was not. Nevertheless, aug-cc-pVXZ basis sets feature ill-conditioned overlap matrices for large molecules, making the aug-MOLOPT basis sets numerically more robust for larger molecules . We report in App. C an example of a calculation of these orbital basis sets on the 9,10-Dihydroanthracene molecule, showing the variation of the error for all the test cases of Fig. 3 with respect to the basis set size.
3 RI basis sets
3.1 Auxiliary RI basis set generation
The resolution-of-the-identity (RI) technique is widely used to reduce the computational cost of quantum chemical methods Vahtras1993. In RI, four-center integrals